A BODY IS PROJECTED HORIZONTALY FROM THE SURFACE OF THE EARTH(RADIUS=R) WITH A VELOCITY EQUAL TO  N TIMES THE ESCAPE VELOCITY. NEGLECT ROTATIONAL EFFECT OF THE EARTH. THE MAXIMUM HEIGHT ATTAINEDBY THE BODY FROM THE EARTH SURFACE IS R/2. THE VALUE OF "N" IS---------------

Grade:12

1 Answers

vikas askiitian expert
13 years ago

H/(R+H) = N2

R/2/(3R/2) = N2

 N = (1/3)1/2